16 years from now ginas age will be twice her age 12 years ago. find the present
Answer by ptfile(81)

16 years from now Gina's age will be twice her age 12 years ago.
G+16=2(G-12)
G+16=2G-24
40=G
Gina's current age is 40 years old.